### v1

# Frozen Gaussian Grid-point Correction For Semi-classical Schrödinger Equation

Abstract
We propose an efficient reconstruction algorithm named the frozen Gaussian grid-point correction (FGGC) for computing the Schrödinger equation in the semi-classical regime using the frozen Gaussian approximation (FGA). The FGA has demonstrated its superior efficiency in dealing with semi-classical problems and high-frequency wave propagations. However, reconstructing the wave function from a large number of Gaussian wave-packets is typically computationally intensive. This difficulty arises because these wave-packets propagate along the FGA trajectories to non-grid positions, making the application of the fast Fourier transform infeasible. In this work, we introduce the concept of “on-grid correction” and derive the formulas for the least squares approximation of Gaussian wave-packets, and also provide a detailed process of the FGGC algorithm. Furthermore, we rigorously prove that the error introduced by the least squares approximation on each Gaussian wave-packet is independent of the semi-classical parameter $\varepsilon$. Numerical experiments show that the FGGC algorithm can significantly improve reconstruction efficiency while introducing only negligible error, making it a powerful tool for solving the semi-classical Schrödinger equation, especially in applications requiring both accuracy and efficiency.
Keywords: Schrödinger equation, semi-classical regime, frozen Gaussian approximation, least square approximation
1 Introduction
The Schrödinger equation resides at the heart of quantum mechanics, serving as the fundamental framework for describing how quantum systems evolve over time [8]. Research on the Schrödinger equation is instrumental in unveiling the fundamental properties of matter and driving advancements in fields such as materials science, condensed matter physics, and quantum information science [18]. In this paper, we shall consider the numerical methods for the semi-classical Schrödinger equation
$\mathrm{i}\varepsilon\frac{\partial u}{\partial t}=-\frac{\varepsilon^{2}}{2}\Delta u+V(\bm{x})u,\ \ \ \ (t,\bm{x})\in\mathbb{R}\times\mathbb{R}^{d},$ (1.1)
with WKB initial condition
$u_{0}(\bm{x})=\sqrt{n_{0}(\bm{x})}\operatorname{exp}\left(\frac{\mathrm{i}}{\varepsilon}S_{0}(\bm{x})\right).$ (1.2)
Here, $\varepsilon$ denotes the semi-classical parameter that describes the microscopic and macroscopic scale ratio and $0<\varepsilon\ll 1$ in the semi-classical regime, $d$ is the dimensionality, $\mathrm{i}=\sqrt{-1}$ is the imaginary unit, $u(t,\bm{x})$ represents the complex-valued wave function, $V(\bm{x})$ is a given real-valued potential function, $\Delta=\Delta_{\bm{x}}$ is the Laplace operator in $\bm{x}$, $n_{0}(\bm{x})$ and $S_{0}(\bm{x})$ are smooth, real-valued functions and independent of $\varepsilon$, and $n_{0}(\bm{x})$ decays sufficiently fast when $\left|\bm{x}\right|\to\infty$.
The properties of the Schrödinger equation is widely studied in both theoretical and numerical aspects. It is well known that the wave function $u(t,\bm{x})$ features oscillations of wave length $\varepsilon$ in space and time and does not converge strongly when $\varepsilon\to 0^{+}$. However, the weak convergence of the wave function is not sufficient for passing to the limit in the macroscopic densities. Therefore, theoretical analysis for the semi-classical limit is a mathematically challenging task. Since the 1990s, significant progress has been made in the study of the semi-classical limit of the Schrödinger equation. Utilizing microlocal analysis tools such as H-measures [26] and Wigner measures [19, 9, 28, 1], researchers have rigorously proved that the semi-classical limit of the Schrödinger equation for a series of complex physical systems is described by classical dynamical equations in phase space. These results illuminate the connection and transition between quantum mechanics and classical mechanics.
Meanwhile, the numerical solution of the Schrödinger equation also requires huge computational resources to achieve accurate physical observables when $\varepsilon$ is small. The challenge becomes even more pronounced when accurate wave functions are of interest [13]. The existing numerical algorithms for solving the Schrödinger equation can be broadly classified into two categories: direct discretization methods and asymptotic methods. Direct discretization methods, such as the finite difference method [6, 7] and the time-splitting spectral method (TSSP) [2], have been developed over several decades, forming a robust class of algorithms with high accuracy. However, these methods often suffer from the drawbacks of low efficiency and high mesh cost: the time-splitting spectral method requires $\Delta t=O(\varepsilon)$ and $\Delta x=O(\varepsilon)$ in order to guarantee an accurate approximation of the wave function, and the finite-difference method even requires $\Delta t=o(\varepsilon)$ and $\Delta x=o(\varepsilon)$. On the other hand, asymptotic methods such as the WKB approximation [24], the Gaussian beam method [14], and the frozen Gaussian approximation (FGA) [3, 20], may cost relatively less computational resources compared to direct discretization methods in the semi-classical regime. However, the WKB solution will become unreliable after the formation of caustics, and the Gaussian beam methods also lose accuracy when the beam evolves and propagates over time, causing its width to increase.
The frozen Gaussian approximation was originally proposed in quantum chemistry to simulate the Schrödinger equation [11, 15, 16, 25, 17], using fixed-width Gaussian functions to approximate the wave function. Recently, FGA theory has been extended and applied to a variety of systems and equations successfully, such as linear strictly hyperbolic systems [21], non-strictly hyperbolic systems including elastic wave equations [10] and relativistic Dirac equations in the semi-classical regime [4, 5], non-adiabatic dynamics in surface hopping problems [12, 23, 22], and open quantum system [27]. These investigations show the adaptability and effectiveness of the FGAs in solving intricate problems in quantum mechanics and high-frequency wave phenomena.
In this paper, we aim to design an efficient reconstruction algorithm for the FGA. Practical experience shows that one needs to consume much more computational resources on the wave reconstruction step compared to the initial decomposition step. The initial decomposition can be performed on grid-points and utilize the fast Fourier transform (FFT) to accelerate the calculation of the numerical integral. However, this approach is not directly applicable to the wave reconstruction step, as the grid-points evolve over time into off-grid points, complicating the reconstruction process. To enhance reconstruction efficiency, we propose an idea to decompose the off-grid Gaussian wave-packets into several on-grid Gaussian wave-packets, which is easy to implement using the least squares approximation and features remarkable efficiency levels. This algorithm is named the frozen Gaussian grid-point correction (FGGC). Then error estimate and numerical experiments are conducted, in which we rigorously prove that the error of the decomposition on a single Gaussian wave-packet is independent of $\varepsilon$, and the overall error of the FGGC algorithm is also independent of $\varepsilon$ numerically. A multi-step solver is also proposed to further improve efficiency for long-time simulation. The results show that, using a basic “on-grid correction” idea, the FGGC algorithm is able to improve the efficiency of the reconstruction step by tens of times, with only a small error cost, which is usually negligible and acceptable.
The remainder of this paper is organized as follows: In section 2, we begin with a brief review of the FGA formulation and then consider the least squares approximation (LSA) for Gaussian wave-packets. Building upon this LSA problem, we present the basic procedure of the frozen Gaussian grid-point correction algorithm. Subsequently, in section 3, we perform an error estimate for the FGGC algorithm. Additionally, we propose improved versions of the FGGC, which further reduces the computational cost. A series of numerical experiments are conducted in section 4, including cases in one to three dimensions. These results demonstrate that the FGGC algorithm significantly improves computational efficiency during wave reconstruction while introducing only a minor error. Moreover, the FGGC algorithm can further enhance the computing performance for long-time evolution problems by employing the multi-step version solver. Section 5 concludes this paper.
2 Frozen Gaussian Approximation and Frozen Gaussian Grid-point Correction
In this section, we briefly recap the frozen Gaussian approximation to the semi-classical Schrödinger equation and introduce the FGGC algorithm employing least squares approximations for wave-packets.
2.1 The frozen Gaussian approximation for semi-classical Schrödinger equation
The frozen Gaussian approximation is an asymptotic approximation to equation eq. 1.1, which gives an integral representation to the solution of wave function:
$u_{\text{FGA}}(t,\bm{x})=\frac{1}{(2\pi\varepsilon)^{\frac{3d}{2}}}\int_{\mathbb{R}^{2d}}A(t,\bm{q},\bm{p})e^{\frac{\mathrm{i}}{\varepsilon}\Phi(t,\bm{x},\bm{q},\bm{p})}\ \mathrm{d}\bm{p}\ \mathrm{d}\bm{q},$ (2.1)
where the phase function is defined as
$\Phi(t,\bm{x},\bm{q},\bm{p})=S(t,\bm{q},\bm{p})+\bm{P}(t,\bm{q},\bm{p})\cdot(\bm{x}-\bm{Q}(t,\bm{q},\bm{p}))+\frac{\mathrm{i}}{2}|\bm{x}-\bm{Q}(t,\bm{q},\bm{p})|^{2}.$ (2.2)
In this integral representation, the position center $\bm{Q}$, momentum center $\bm{P}$, amplitude $A$ and action $S$ are time-dependent variables satisfying an ODE system:
$\frac{\mathrm{d}\bm{Q}}{\mathrm{d}t}$ $=\bm{P},$ $\bm{Q}(0,\bm{q},\bm{p})=\bm{q},$ (2.3)
$\frac{\mathrm{d}\bm{P}}{\mathrm{d}t}$ $=-\nabla V(\bm{Q}),$ $\bm{P}(0,\bm{q},\bm{p})=\bm{p},$ (2.4)
$\frac{\mathrm{d}S}{\mathrm{d}t}$ $=\frac{1}{2}|\bm{P}|^{2}-V(\bm{Q}),$ $S(0,\bm{q},\bm{p})=0,$ (2.5)
$\frac{\mathrm{d}A}{\mathrm{d}t}$ $=\frac{A}{2}\operatorname{tr}\left(\bm{Z}^{-1}\cdot\frac{\mathrm{d}\bm{Z}}{\mathrm{d}t}\right),$ $A(0,\bm{q},\bm{p})=2^{\frac{d}{2}}\widehat{u_{0}}(\bm{q},\bm{p}),$ (2.6)
where
$\widehat{u}_{0}(\bm{q},\bm{p})=\int_{\mathbb{R}^{d}}u_{0}(\bm{y})e^{\frac{\mathrm{i}}{\varepsilon}\big{(}-\bm{p}\cdot(\bm{y}-\bm{q})+\frac{\mathrm{i}}{2}|\bm{y}-\bm{q}|^{2}\big{)}}\mathrm{d}\bm{y}.$ (2.7)
In the above ODE system, we have used the shorthand notations:
$\begin{split}\partial_{\bm{z}}=\partial_{\bm{q}}-\mathrm{i}\partial_{\bm{p}}\ \ \ \operatorname{and}\ \ \ \bm{Z}=\partial_{\bm{z}}(\bm{Q}+\mathrm{i}\bm{P}),\end{split}$ (2.8)
where $\partial_{\bm{z}}\bm{Q}$,$\partial_{\bm{z}}\bm{P}$ and $\bm{Z}$ are understood as matrices. A detailed derivation for formulas eq. 2.3$\simeq$. 2.7 can be found in [20]. The following proposition shows that the FGA ansatz given by eq. 2.1 is exact for the initial condition eq. 2.7.
Proposition 1.
Let $u_{0}(\bm{x})\in L^{2}(\mathbb{R}^{d})$ be the initial condition of the Schrödinger equation, then we have
$u_{0}(\bm{x})=\frac{1}{(2\pi\varepsilon)^{\frac{3d}{2}}}\int_{\mathbb{R}^{2d}}A(0,\bm{q},\bm{p})e^{\frac{\mathrm{i}}{\varepsilon}\Phi(0,\bm{x},\bm{q},\bm{p})}\mathrm{d}\bm{q}\ \mathrm{d}\bm{p},$ (2.9)
where the initial condition is given by eq. 2.3$\simeq$. 2.7.
The FGA approximates the Schrödinger equation by three steps: First, decompose the initial waves into several Gaussian wave-packets in the phase space to determine $A(0,\bm{q},\bm{p})$ by eq. 2.7 for each $(\bm{q},\bm{p})$. Second, propagate the wave-packets along the Hamiltonian flow by eq. 2.3$\simeq$. 2.6. Third, reconstruct the wave function using the wave-packets after propagation.
In practical algorithms, the FGA method requires the calculation of a series of discrite summation to approximate the integrals in initial decomposition and wave reconstruction. For initial decomposition, FFT can be utilized to accelerate the calculation. However for the wave reconstruction, the process can be approximated by the following discrete form
$u(t,\bm{x})$ $\approx\ \frac{(\Delta q\cdot\Delta p)^{d}}{(2\pi\varepsilon)^{\frac{3d}{2}}}\sum_{j=1}^{M}A_{j}\operatorname{exp}\left(\frac{\mathrm{i}}{\varepsilon}\left(S_{j}+\bm{P}_{j}\cdot\left(\bm{x}-\bm{Q}_{j}\right)\right)-\frac{1}{2\varepsilon}\left(\bm{x}-\bm{Q}_{j}\right)^{2}\right),$ (2.10)
in which FFT is no longer feasible. This is due to the fact that the wave-packets generally do not align with the grid points of the phase space. Consequently, the computational time for wave reconstruction significantly exceeds that of the initial decomposition, and this computational time escalates rapidly with an increase in dimensionality $d$ or a decrease in the semi-classical parameter $\varepsilon$.
2.2 Least squares approximation problem and formulas for Gaussian wave-packets
As mentioned earlier, summing up Gaussian wave-packets centered at off-grid points demands a significant amount of time. To address this issue, we propose a method for decomposing an off-grid Gaussian wave-packet into several on-grid Gaussian wave-packets. A natural idea is to utilize the least squares approximation (LSA), which offers an “optimal” approximation in the $L^{2}$ sense.
Generally, consider the least squares approximation of a complex-valued function $\phi(x)$ in a function space $V$ spanned by a set of basis functions $\left\{\psi^{(k)}\right\}_{k=1}^{n}$. We can express the approximation by
$\begin{split}\phi^{*}\approx\sum_{k=1}^{n}c^{(k)}\psi^{(k)}.\end{split}$ (2.11)
The coefficients are determined by the normal equations:
$\begin{split}\sum_{k=1}^{n}A_{jk}c^{(k)}=f_{j},\ \ j=1,2,\dots,n,\end{split}$ (2.12)
where
$A_{jk}=\left<\psi^{(j)},\psi^{(k)}\right>\text{ and }f_{j}=\left<\psi^{(j)},\phi\right>,$ (2.13)
[…]
$\begin{split}\langle g,h\rangle=\int_{\mathbb{R}^{d}}g(\bm{x})\cdot h(\bm{x})^{\dagger}\ \mathrm{d}\bm{x}.\end{split}$ (2.14)
In the FGA ansatz, the given function and the approximating functions are in the forms of fixed-width Gaussian wave-packet. For instance:
$\begin{split}\phi(\bm{x})=G\left(\bm{x};\bm{Q}_{0},\bm{P}_{0},\varepsilon\right)\text{ and }\psi^{(k)}(\bm{x})=G\left(\bm{x};\bm{Q}_{0}^{(k)},\bm{P}_{0}^{(k)},\varepsilon\right),\end{split}$ (2.15)
in which $\bm{Q}_{0}^{(k)}$ and $\bm{P}_{0}^{(k)}$ should be the grid-points near $\bm{Q}_{0}$ and $\bm{P}_{0}$ respectively. Here, $G(\bm{x};\bm{Q},\bm{P},\varepsilon)$ denotes the Gaussian wave-packet
$G(\bm{x};\bm{Q},\bm{P},\varepsilon)=e^{\frac{\mathrm{i}}{\varepsilon}\bm{P}\cdot(\bm{x}-\bm{Q})-\frac{1}{2\varepsilon}|\bm{x}-\bm{Q}|^{2}}.$ (2.16)
For simplicity, we may omit $\varepsilon$ and write $G(\bm{x};\bm{Q},\bm{P},\varepsilon)$ as $G(\bm{x};\bm{Q},\bm{P})$ when $\varepsilon$ is fixed and not of concern.
Essentially, these inner products are interactions between two Gaussian wave-packets and are straightforward to compute. For instance, if $g(\bm{x})=G(\bm{x};\bar{\bm{Q}},\bar{\bm{P}})$ and $h(\bm{x})=(\bm{x};\widetilde{\bm{Q}},\widetilde{\bm{P}})$, then
$g,h\right>&=\int_{\mathbb{R}^{d}}g(\bm{x})\,h(\bm{x})^{\dagger}\ \mathrm{d}\bm{x}\\ &=(\pi\varepsilon)^{\frac{d}{2}}\operatorname{exp}\left(-\frac{1}{4\varepsilon}\left|\widetilde{\bm{Q}}-\bar{\bm{Q}}\right|^{2}-\frac{1}{4\varepsilon}\left|\widetilde{\bm{P}}-\bar{\bm{P}}\right|^{2}+\frac{\mathrm{i}}{2\varepsilon}(\widetilde{\bm{Q}}-\bar{\bm{Q}})\cdot(\widetilde{\bm{P}}+\bar{\bm{P}})\right).\end{split}$ (2.17)
So, substitute $\widetilde{\bm{Q}},\bar{\bm{Q}},\widetilde{\bm{P}},\bar{\bm{P}}$ with $\bm{Q}^{(j)},\bm{Q}^{(k)},\bm{P}^{(j)},\bm{P}^{(k)}$, one can calculate all the entries of the coefficient matrix $\bm{A}$ and right-hand side vector $\bm{f}$ of equation eq. 2.12. For example, for $d=1$, two nearby points can be chosen for each component as follows:
Figure 1: Splitting one off-grid wave-packet into four on-grid wave-packets
$Q^{(1)}$ $=\left(\left[\frac{Q}{\Delta q}\right]\right)\cdot\Delta q,$ $P^{(1)}=\left(\left[\frac{P}{\Delta p}\right]\right)\cdot\Delta p,$ (2.18)
$Q^{(2)}$ $=\left(\left[\frac{Q}{\Delta q}\right]+1\right)\cdot\Delta q,$ $P^{(2)}=\left(\left[\frac{P}{\Delta p}\right]\right)\cdot\Delta p,$
$Q^{(3)}$ $=\left(\left[\frac{Q}{\Delta q}\right]\right)\cdot\Delta q,$ $P^{(3)}=\left(\left[\frac{P}{\Delta p}\right]+1\right)\cdot\Delta p,$
$Q^{(4)}$ $=\left(\left[\frac{Q}{\Delta q}\right]+1\right)\cdot\Delta q,$ $P^{(4)}=\left(\left[\frac{P}{\Delta p}\right]+1\right)\cdot\Delta p.$
[…]
Figure 2: The LSA error and coefficients, in the case of $d=1$, $4$ neighbours, $\varepsilon=1/64$, $\Delta q=1/16$, $\Delta p=\pi/32$, $0\leq Q_{0}<\Delta q$, $0\leq P_{0}<\Delta p$
2.3 The frozen Gaussian grid-point correction algorithm
Suppose that for any $(\bm{q},\bm{p})$ in the phase space, $A(0,\bm{q},\bm{p})$ has already been calculated by eq. 2.7. By evolving the ODE system eq. 2.3 $\sim$ eq. 2.7 up to time $t$, one can reconstruct the wave function as
$u(t,\bm{x})$ $\approx\ \frac{(\Delta q\cdot\Delta p)^{d}}{(2\pi\varepsilon)^{\frac{3d}{2}}}\sum_{j=1}^{M}A_{j}\cdot e^{\frac{\mathrm{i}}{\varepsilon}S_{j}}\cdot G\left(\bm{x};\bm{Q}_{j},\bm{P}_{j}\right).$ (2.19)
Here, $\Delta q$ and $\Delta p$ denote the discrete step size for numerical integration, $M$ denotes the number of wave-packets included in the calculation. For simplicity, we omit the explicit time dependence such as $A_{j}=A_{j}(t)$ and adopt the notations $G_{j}(\bm{x})=G\left(\bm{x};\bm{Q}_{j},\bm{P}_{j}\right)$ in the following text. For any $j$, using the formulas introduced previously, we can perform the least squares approximation for each $G_{j}(\bm{x})$. We first introduce some notations and definitions for choosing a set of neighbours.
Definition 1.
Assume that the semi-classical parameter $\varepsilon$ and the discrete mesh step $\Delta q$ and $\Delta p$ are given. Then, choose $2n$ sequences of shift integers $\bm{\delta q}^{(k)}=\left(\delta q^{(k)}_{l}\right)_{l=1}^{d}$ and $\bm{\delta p}^{(k)}=\left(\delta p^{(k)}_{l}\right)_{l=1}^{d}$ for $k=1,2,\cdots,n$, and define the neighbour points map $\bm{\mathcal{N}}(\bm{Q},\bm{P};k)=\left(\bm{Q}^{(k)},\bm{P}^{(k)}\right)$ by:
$\bm{Q}^{(k)}=$ $\operatorname{floor}\left(\frac{\bm{Q}}{\Delta q}\right)\Delta q+\bm{\delta q}^{(k)}\Delta q,$ (2.20)
$\bm{P}^{(k)}=$ $\operatorname{floor}\left(\frac{\bm{P}}{\Delta p}\right)\Delta p+\bm{\delta p}^{(k)}\Delta p,$
where $\operatorname{floor}(\cdot)$ denotes the floor function.
Therefore, the neighbour points map eq. 2.20 can be interpreted as a “neighbour strategy”: selecting a set of $\delta q_{l}^{(k)}$ and $\delta p_{l}^{(k)}$ generates $n$ distinct on-grid neighbour wave-packets for the target wave-packet. Now we choose a set of neighbours for each specific wave-packet $G_{j}$, where the $k$-th neighbour can be represented as
$G_{j}^{(k)}(\bm{x}):=G(\bm{x};\bm{\mathcal{N}}(\bm{Q}_{j},\bm{P}_{j};k)).$ (2.21)
Then the least squares approximation
$G_{j}(\bm{x})\approx\sum_{k=1}^{n}c_{j}^{(k)}\cdot G_{j}^{(k)}(\bm{x}),$ (2.22)
can be obtained by solving the normal equations eq. 2.12 with $\phi=G_{j}$ and $\psi^{(k)}=G_{j}^{(k)}$. Substituting eq. 2.22 into eq. 2.19 yields
$u(t,\bm{x})\approx\ \frac{(\Delta q\cdot\Delta p)^{d}}{(2\pi\varepsilon)^{\frac{3d}{2}}}\sum_{j=1}^{M}A_{j}e^{\frac{\mathrm{i}}{\varepsilon}S_{j}}\sum_{k=1}^{n}c_{j}^{(k)}G_{j}^{(k)}(\bm{x}).$ (2.23)
Note that the parameters of $G_{j}^{(k)}$, namely $\bm{Q}^{(k)}_{j}$ and $\bm{P}^{(k)}_{j}$, are all on-grid. Thus, we can combine Gaussian wave-packets sharing the same $\bm{Q}_{j}^{(k)}$ and $\bm{P}_{j}^{(k)}$:
$u(t,\bm{x})\approx\frac{(\Delta q\cdot\Delta p)^{d}}{(2\pi\varepsilon)^{\frac{3d}{2}}}\sum_{(\bm{q},\bm{p})\operatorname{\ on\ grid}}\widetilde{A}(\bm{q},\bm{p})\cdot G(\bm{x};\bm{q},\bm{p}),$ (2.24)
in which
$\widetilde{A}(\bm{q},\bm{p})=\sum_{\begin{subarray}{c}\bm{Q}_{j}^{(k)}=\bm{q}\\ \bm{P}_{j}^{(k)}=\bm{p}\end{subarray}}A_{j}\cdot e^{\frac{\mathrm{i}}{\varepsilon}S_{j}}\cdot c_{j}^{(k)}.$ (2.25)
Now the summation eq. 2.24 has such a form so that FFT can be utilized directly in order to enhance efficiency. In summary, we present the pseudo-code for the frozen Gaussian grid-point correction (FGGC) in algorithm 1
Algorithm 1 Frozen Gaussian grid-point correction (FGGC) - A basic version
  Step 0: Preparation
  Discrete the meshes, take $\Delta x=O(\varepsilon),\Delta q=O(\sqrt{\varepsilon}),\Delta p=O(\sqrt{\varepsilon})$.
  Step 1: Initial decomposition
  For each $\bm{q}$, compute $A(0,\bm{q},\bm{p})$ using FFT algorithm by eq. 2.7.
  Step 2: Time evolution
  For each wave-packet, evolve the ODE system eq. 2.3$\simeq$. 2.6 up to time $t$, and obtain the FGA variables $A,S,\bm{Q},\bm{P}$.
  Step 3: Least squares approximation
  Determine the neighbour strategy as in eq. 2.20 and eq. 2.21. For each wave-packet, solve the normal equations eq. 2.12 and obtain the LSA coefficients.
  Step 4: Wave reconstruction
  Compute the wave function based on eq. 2.24 and eq. 2.25.
Remark 1: The transform in eq. 2.7 consists of a sequence of cut-off FFT. Consequently, one could take $\Delta p=O(\sqrt{\varepsilon})$ as the step size of the phase space mesh. In fact, through stationary phase approximation, one gets that $A(0,\bm{q},\bm{p})$ is localized around the submanifold $\bm{p}=\nabla_{\bm{q}}S_{0}(\bm{q})$, which inspires one to put the mesh grids of $\bm{p}$ around $\nabla_{\bm{q}}S_{0}(\bm{q})$ to perform the initial decomposition. Additionally, taking $\Delta q=O(\sqrt{\varepsilon})$ also ensures the accuracy of the wave function.
[…]
Remark 3: The FGGC algorithm is highly parallelizable. For Step 1, parallel computation can be performed in the $\bm{q}$ direction. Specifically, to obtain $A(0,\bm{q},\bm{p})$, the $\mathrm{d}\bm{y}$ integral eq. 2.7 can be computed separately (using FFT) for each $\bm{q}$. For Step 2 and Step 3, the evolution and least squares approximation of each wave-packet can also be carried out in parallel. For Step 4, the $\mathrm{d}\bm{p}$ integral eq. 2.24 can be computed separately (using FFT) for each $\bm{q}$, and then sum them up to obtain $u(t,\bm{x})$.
3 Error Estimate and Efficiency Analysis for FGGC
In this section, we conduct an error analysis of the FGGC algorithm. Obviously, the $L^{2}$ error of the FGGC wave function $u_{\text{fggc}}$ with respect to the exact solution $u_{\text{exact}}$ can be divided into two parts:
$\left\|u_{\text{exact}}-u_{\text{fggc}}\right\|_{L^{2}}\leq\left\|u_{\text{exact}}-u_{\text{fga}}\right\|_{L^{2}}+\left\|u_{\text{fga}}-u_{\text{fggc}}\right\|_{L^{2}}.$ (3.1)
The following theorem shows that the asymptotic error of the FGA ansatz, i.e. the first term on the right-hand side of eq. 3.1 is $O(\varepsilon)$:
Theorem 1.
[20] Assume the potential function $V(\bm{x})\in C^{\infty}(\mathbb{R}^{\textup{d}})$. Let $u_{\textup{exact}}$ be the solution of the semi-classical Schrödinger equation eq. 1.1eq. 1.2, and $u_{\textup{FGA}}$ be the FGA representation as eq. 2.1. For any $t>0$, we have
$\left\|u_{\textup{exact}}(t,\cdot)-u_{\textup{fga}}(t,\cdot)\right\|_{L^{2}}\leq C_{1}(t,d)\left\|u_{0}\right\|_{L^{2}}\cdot\varepsilon,$ (3.2)
where $C_{1}(t,d)$ is a parameter independent of the semi-classical constant $\varepsilon$.
For the second term, we simply rewrite the difference as
$u_{\text{fga}}(t,\cdot)-u_{\text{fggc}}(t,\cdot)=\frac{(\Delta q\cdot\Delta p)^{d}}{(2\pi\varepsilon)^{\frac{3d}{2}}}\sum_{j=1}^{M}A_{j}\cdot\left(G_{j}(\bm{\cdot})-\sum_{k=1}^{n}c_{j}^{(k)}G_{j}^{(k)}(\bm{\cdot})\right).$ (3.3)
A direct estimate of this expression is challenging. However, we can try to estimate only the difference between a wave-packet and its least squares approximation.
3.1 Error estimate for LSA problem and numerical comparison on different strategies
We denote the relative $L^{2}$ error of the least squares approximation for a single Gaussian wave-packet as
$E_{\text{LSA}}:=$ $\frac{\left\|G\left(\bm{x};\bm{Q},\bm{P}\right)-\sum_{k=1}^{n}c^{(k)}G\left(\bm{x};\bm{\mathcal{N}}(\bm{Q},\bm{P};k)\right)\right\|_{L^{2}}}{\left\|G(\bm{x};\bm{Q},\bm{P})\right\|_{L^{2}}}$ (3.4)
$=$ $(\pi\varepsilon)^{-\frac{d}{4}}\left\|G(\bm{x};\bm{Q},\bm{P})-\sum_{k=1}^{n}c^{(k)}G\left(\bm{x};\bm{\mathcal{N}}(\bm{Q},\bm{P};k)\right)\right\|_{L^{2}}.$
Intuitively, $E_{\text{LSA}}=E_{\text{LSA}}(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p)$ depends on the following variables:
(1) The semi-classical parameter $\varepsilon$;
(2) The coordinate of the target $(\bm{Q},\bm{P})$;
(3) The neighbour strategy, including the number of the neighbours $n$ and the shift integers $\delta q_{l}^{(k)}$ with $\delta p_{l}^{(k)}$ as in eq. 2.20;
(4) The mesh step size $\Delta q$ and $\Delta p$.
We adopt the following conventions for the “mesh strategy” and the “neighbour strategy”: The mesh strategy is considered fixed, if
$\Delta q=C_{q}\sqrt{\varepsilon},\ \Delta p=C_{p}\sqrt{\varepsilon},$ (3.5)
for some constants $C_{q}$ and $C_{p}$. The neighbour strategy is considered fixed, if the number of neighbours $n$ and the sequences of shift integers $\bm{\delta q}^{(k)}$ and $\bm{\delta p}^{(k)}$ for and $k=1,2,\cdots,n$ are all given. To study the properties of $E_{\text{LSA}}(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p)$, we introduce more notations.
Notations. Denote
$\operatorname{int}\bm{Q}=\operatorname{floor}\left(\frac{\bm{Q}}{\Delta q}\right)\cdot\Delta q,\ \operatorname{int}\bm{P}=\operatorname{floor}\left(\frac{\bm{P}}{\Delta p}\right)\cdot\Delta p,$ (3.6)
and
$\operatorname{frac}\bm{Q}=\bm{Q}-\operatorname{int}\bm{Q},\ \operatorname{frac}\bm{P}=\bm{P}-\operatorname{int}\bm{P},$ (3.7)
where $\operatorname{floor}\left(\cdot\right)$ represents the floor function. Therefore, the phase space coordinate $(\bm{Q},\bm{P})$ is decomposed into “integer part” and “fractional part”, where the “fractional part” always falls within the range of $[0,\Delta q)^{d}\times[0,\Delta p)^{d}$. We refer to this range as the “unit hypercube” in the following text.
Applying transformations to eq. 3.4 leads to the following theorem:
Theorem 2.
Suppose the mesh strategy and the neighbour strategy are fixed. Consider the relative LSA error $E_{\text{LSA}}(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p)$, we have
$E_{\text{LSA}}\left(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p\right)=E_{\text{LSA}}\left(\varepsilon,\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P},\Delta q,\Delta p\right).$ (3.8)
Proof.
Denote $\left\{c^{(k)}\right\}_{k=1}^{n}$ as the LSA coefficients with respect to $\left(\bm{Q},\bm{P}\right)$ and $\left\{\widetilde{c}^{(k)}\right\}_{k=1}^{n}$ as that of $\left(\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P}\right)$. Recall that $\left\{c^{(k)}\right\}_{k=1}^{n}$ satisfy the normal equations eq. 2.12 and eq. 2.13, while $\left\{\widetilde{c}^{(k)}\right\}_{k=1}^{n}$ satisfy
$\sum_{k=1}^{n}\widetilde{A}_{jk}\widetilde{c}^{(k)}=\widetilde{f}_{j}\ ,j=1,2,\cdots,n,$ (3.9)
$\widetilde{A}_{jk}=$ $G(\bm{x};\bm{\mathcal{N}}(\bm{0},\bm{0};j)),G(\bm{x};\bm{\mathcal{N}}(\bm{0},\bm{0};k))\right>,$
$\widetilde{f}_{j}=$ $G(\bm{x};\bm{\mathcal{N}}(\bm{0},\bm{0};j)),G(\bm{x};\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P})\right>.$
Using eq. 2.17, one gets
$\widetilde{A}_{jk}=A_{jk}\operatorname{exp}\left(-\frac{\mathrm{i}}{\varepsilon}\left(\bm{Q}^{(k)}-\bm{Q}^{(j)}\right)\cdot\operatorname{int}\bm{P}\right),$ (3.10)
and
$\widetilde{f}_{j}=f_{j}\operatorname{exp}\left(-\frac{\mathrm{i}}{\varepsilon}\left(\bm{Q}-\bm{Q}^{(j)}\right)\cdot\operatorname{int}\bm{P}\right).$ (3.11)
So, eq. 3.9 becomes
$\sum_{k=1}^{n}A_{jk}\operatorname{exp}\left(\frac{\mathrm{i}}{\varepsilon}\left(\bm{Q}-\bm{Q}^{(k)}\right)\cdot\operatorname{int}\bm{P}\right)\widetilde{c}^{(k)}=f_{j}\ ,j=1,2,\cdots,n.$ (3.12)
Therefore,
$c^{(k)}=\widetilde{c}^{(k)}\operatorname{exp}\left(\frac{\mathrm{i}}{\varepsilon}\left(\bm{Q}-\bm{Q}^{(k)}\right)\cdot\operatorname{int}\bm{P}\right).$ (3.13)
Next, perform a change of variable $\bm{x}\to\bm{x}+\operatorname{int}\bm{Q}$, which is equivalent to replacing $\bm{Q}$ with $\operatorname{frac}\bm{Q}$, we have
$E_{\text{LSA}}(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p)$ (3.14)
$=$ $(\pi\varepsilon)^{-\frac{d}{4}}\left\|G\left(\bm{x};\operatorname{frac}\bm{Q},\bm{P}\right)-\sum_{k=1}^{n}c^{(k)}G\left(\bm{x};\bm{\mathcal{N}}(\operatorname{frac}\bm{Q},\bm{P};k)\right)\right\|_{L^{2}}$
$=$ $(\pi\varepsilon)^{-\frac{d}{4}}\left\|G(\bm{x};\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P})e^{\frac{\mathrm{i}}{\varepsilon}(\bm{x}-\operatorname{frac}\bm{Q})\cdot\operatorname{int}\bm{P}}-\sum_{k=1}^{n}c^{(k)}G\left(\bm{x};\bm{\mathcal{N}}(\operatorname{frac}\bm{Q},\bm{P};k)\right)\right\|_{L^{2}}$
$=$ $(\pi\varepsilon)^{-\frac{d}{4}}\left\|G(\bm{x};\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P})-\sum_{k=1}^{n}c^{(k)}e^{-\frac{\mathrm{i}}{\varepsilon}(\bm{x}-\operatorname{frac}\bm{Q})\cdot\operatorname{int}\bm{P}}G\left(\bm{x};\bm{\mathcal{N}}(\operatorname{frac}\bm{Q},\bm{P};k)\right)\right\|_{L^{2}}$
$=$ $(\pi\varepsilon)^{-\frac{d}{4}}\left\|G(\bm{x};\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P})-\sum_{k=1}^{n}\widetilde{c}^{(k)}G\left(\bm{x};\bm{\mathcal{N}}(\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P};k)\right)\right\|_{L^{2}}$
$=$ $E_{\text{LSA}}(\varepsilon,\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P},\Delta q,\Delta p).$
Then the statement is proved. ∎
Using this theorem, the domain of the phase space coordinate can be restricted in the “unit hypercube” when considering the supremum of the relative error, i.e.
$\sup_{(\bm{Q},\bm{P})\in\mathbb{R}^{2d}}E_{\text{LSA}}(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p)=$ $\sup_{(\bm{Q},\bm{P})\in\mathbb{R}^{2d}}E_{\text{LSA}}(\varepsilon,\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P},\Delta q,\Delta p)$ (3.15)
$=$ $\sup_{(\bm{Q},\bm{P})\in[0,\Delta q]^{d}\times[0,\Delta p]^{d}}E_{\text{LSA}}(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p).$
Next, we investigate the relationship between $E_{\text{LSA}}$ and the semi-classical parameter $\varepsilon$.
Theorem 3.
Assume that both the mesh strategy and the neighbour strategy are fixed. Then for $\forall\varepsilon_{1},\varepsilon_{2}>0$, we have:
$\sup_{(\bm{Q},\bm{P})\in[0,\Delta q_{1}]^{d}\times[0,\Delta p_{1}]^{d}}E_{\text{LSA}}(\varepsilon_{1},\bm{Q},\bm{P},\Delta q_{1},\Delta p_{1})$ (3.16)
$=$ $\sup_{(\bm{Q},\bm{P})\in[0,\Delta q_{2}]^{d}\times[0,\Delta p_{2}]^{d}}E_{\text{LSA}}(\varepsilon_{2},\bm{Q},\bm{P},\Delta q_{2},\Delta p_{2}),$
where $\Delta q_{1}=C_{q}\sqrt{\varepsilon_{1}}$, $\Delta p_{1}=C_{p}\sqrt{\varepsilon_{1}}$ and $\Delta q_{2}=C_{q}\sqrt{\varepsilon_{2}}$, $\Delta p_{2}=C_{p}\sqrt{\varepsilon_{2}}$.
Proof.
It suffices to prove that
$E_{\text{LSA}}(\varepsilon_{1},\Delta q_{1}\bm{s},\Delta p_{1}\bm{t},\Delta q_{1},\Delta p_{1})=E_{\text{LSA}}(\varepsilon_{2},\Delta q_{2}\bm{s},\Delta p_{2}\bm{t},\Delta q_{2},\Delta p_{2})$ (3.17)
holds for $\forall(\bm{s},\bm{t})\in[0,1]^{d}\times[0,1]^{d}$.
By making a change of variable $\bm{y}=\frac{\bm{x}}{\sqrt{\varepsilon}}$ in the norm eq. 3.4, a differential term $\varepsilon^{\frac{d}{4}}$ would be introduced in the numerator, which cancels with the same term in the denominator:
$E_{\text{LSA}}(\varepsilon_{1},\Delta q_{1}\bm{s},\Delta p_{1}\bm{t},\Delta q_{1},\Delta p_{1})$ (3.18)
$=\pi^{\frac{d}{4}}\left\|G\left(\sqrt{\varepsilon_{1}}\bm{y};\Delta q_{1}\bm{s},\Delta p_{1}\bm{t},\varepsilon_{1}\right)-\sum_{k=1}^{n}c_{1}^{(k)}G\left(\sqrt{\varepsilon_{1}}\bm{y};\bm{\mathcal{N}}(\Delta q_{1}\bm{s},\Delta p_{1}\bm{t};k),\varepsilon_{1}\right)\right\|_{L^{2}},$
where $\bm{c}_{1}$ are the LSA coefficients corresponding to $E_{\text{LSA}}(\varepsilon_{1},\Delta q_{1}\bm{s},\Delta p_{1}\bm{t},\Delta q_{1},\Delta p_{1})$.
By eq. 2.16, it holds that
$G\left(\sqrt{\varepsilon}\bm{y};\bm{Q},\bm{P},\varepsilon\right)=G\left(\bm{y};\frac{\bm{Q}}{\sqrt{\varepsilon}},\frac{\bm{P}}{\sqrt{\varepsilon}},1\right).$ (3.19)
Therefore,
$E_{\text{LSA}}$ $(\varepsilon_{1},\Delta q_{1}\bm{s},\Delta p_{1}\bm{t},\Delta q_{1},\Delta p_{1})$ (3.20)
$=\pi^{\frac{d}{4}}\left\|G\left(\bm{y};C_{q}\bm{s},C_{p}\bm{t},1\right)-\sum_{k=1}^{n}c_{1}^{(k)}G\left(\bm{y};\bm{\mathcal{N}}\left(C_{q}\bm{s},C_{p}\bm{t};k\right),1\right)\right\|_{L^{2}}.$
Similarly,
$E_{\text{LSA}}$ $(\varepsilon_{2},\Delta q_{2}\bm{s},\Delta p_{2}\bm{t},\Delta q_{2},\Delta p_{2})$ (3.21)
$=\pi^{\frac{d}{4}}\left\|G\left(\bm{y};C_{q}\bm{s},C_{p}\bm{t},1\right)-\sum_{k=1}^{n}c_{2}^{(k)}G\left(\bm{y};\bm{\mathcal{N}}\left(C_{q}\bm{s},C_{p}\bm{t};k\right),1\right)\right\|_{L^{2}},$
where $\bm{c}_{2}$ are the LSA coefficients corresponding to $E_{\text{LSA}}(\varepsilon_{2},\Delta q_{2}\bm{s},\Delta q_{2}\bm{t},\Delta q_{2},\Delta p_{2})$.
By comparing eq. 3.20 and eq. 3.21, it suffices to show that $\bm{c}_{1}=\bm{c}_{2}$. Substituting the neighbour representation eq. 2.20 into the normal equations eq. 2.12 and eq. 2.13, one obtains the coefficient matrix of the normal equations for $\bm{c}_{1}$:
$A_{jk}=\operatorname{exp}\bigg{(}$ $-\frac{C_{q}^{2}}{4}\left|\bm{\delta q}^{(j)}-\bm{\delta q}^{(k)}\right|^{2}-\frac{C_{p}^{2}}{4}\left|\bm{\delta p}^{(j)}-\bm{\delta p}^{(k)}\right|^{2}$ (3.22)
$+\frac{\mathrm{i}C_{q}C_{p}}{2}\left(\bm{\delta q}^{(k)}-\bm{\delta q}^{(j)}\right)\cdot\left(\bm{\delta p}^{(j)}+\bm{\delta p}^{(k)}\right)\bigg{)},$
and the right-hand side vector:
$f_{j}=\operatorname{exp}\bigg{(}$ $-\frac{C_{q}^{2}}{4}\left|\bm{\delta q}^{(j)}-\bm{s}\right|^{2}-\frac{C_{p}^{2}}{4}\left|\bm{\delta p}^{(j)}-\bm{t}\right|^{2}$ (3.23)
$+\frac{\mathrm{i}C_{q}C_{p}}{2}\left(\bm{s}-\bm{\delta q}^{(j)}\right)\cdot\left(\bm{\delta p}^{(j)}+\bm{t}\right)\bigg{)}.$
By performing similar derivations, it is obvious that the normal equations corresponding to $\bm{c}_{2}$ have exactly the same expressions as eq. 3.22 and eq. 3.23, since the above two equations do not contain explicit dependence on $\varepsilon_{1}$. Consequently, it follows that $\bm{c}_{1}=\bm{c}_{2}$, which validates eq. 3.17, thus proving the statement. ∎
Once the discrete mesh strategy and the neighbour strategy are determined, denote the supremum of the relative LSA error $E_{\text{LSA}}$ over all values of $\varepsilon$ and all wave-packets as
$M_{\text{LSA}}:=$ $\sup_{\begin{subarray}{c}\varepsilon>0\\ (\bm{Q},\bm{P})\in\mathbb{R}^{2d}\end{subarray}}E_{\text{LSA}}\left(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p\right),$ (3.24)
where $\Delta q=C_{q}\sqrt{\varepsilon}$ and $\Delta p=C_{p}\sqrt{\varepsilon}$. Then, theorem 2 and theorem 3 bring us closer to achieving our goal of determining $M_{\text{LSA}}$:
$M_{\text{LSA}}=$ $\sup_{\begin{subarray}{c}\varepsilon>0\\ (\bm{Q},\bm{P})\in[0,\Delta q]^{d}\times[0,\Delta p]^{d}\end{subarray}}E_{\text{LSA}}\left(\varepsilon,\bm{Q},\bm{P},\Delta q,\Delta p\right)$ (3.25)
$=$ $\sup_{(\bm{Q},\bm{P})\in[0,\Delta q]^{d}\times[0,\Delta p]^{d}}E_{\text{LSA}}\left(\varepsilon_{0},\bm{Q},\bm{P},\Delta q,\Delta p\right)$
where $\varepsilon_{0}>0$ can be chosen arbitrarily. It is clear that $M_{\text{LSA}}$ is bounded since the supremum is taken over a compact set with respect to $(\bm{Q},\bm{P})$, and depends only on the mesh strategy and the neighbour strategy. At this point, we can assert that the least squares approximation error for a single wave-packet is independent of $\varepsilon$ under the determined mesh strategy and the neighbour strategy.
We then proceed to investigate the numerical behavior of $M_{\text{LSA}}$ in various mesh strategies and neighbour strategies. Take $d=1$, $\varepsilon=1/64$, $\Delta q=2\sqrt{\varepsilon}$, $\sqrt{\varepsilon}$, $\sqrt{\varepsilon}/2$, $\sqrt{\varepsilon}/4$, $\Delta p=\pi\sqrt{\varepsilon}/2$, $\pi\sqrt{\varepsilon}/4$, $\pi\sqrt{\varepsilon}/8$, $\pi\sqrt{\varepsilon}/16$, and several different neighbour strategies as in fig. 3. We will briefly refer to these neighbour strategies as Q2P2, Q2P4, Q4P2, Q4P4. We compute eq. 3.25 numerically for different strategies, and the results are shown in table 4 $\sim$ table 4. It can be observed that both decreasing the step size and increasing the number of neighbours contribute to a reduction in the LSA error, which is consistent with intuitive expectations. When comparing different strategies, we suggest choosing the mesh strategy $\Delta q=\sqrt{\varepsilon}/2,\Delta p=\pi\sqrt{\varepsilon}/8$ and neighbour strategy Q4P2. This is because:
(1) The mesh strategy ensures the accuracy of the FGA;
(2) The value of $M_{\textup{LSA}}$ is sufficiently small under the neighbour strategy;
(3) Decreasing the step size of discretization would result in a greater time cost, since it would increase the number of wave-packets that need to be evolved over time.
The above results can be readily extended to higher-dimensional cases.
Figure 3: Four commonly used neighbour strategies in one dimension
$dp$ $dq$ $2\sqrt{\varepsilon}$ $\sqrt{\varepsilon}$ $\sqrt{\varepsilon}/2$ $\sqrt{\varepsilon}/4$
$\pi\sqrt{\varepsilon}/2$ 1.540E-01 4.815E-02 2.672E-02 2.112E-02
$\pi\sqrt{\varepsilon}/4$ 7.831E-02 1.072E-02 3.141E-03 1.699E-03
$\pi\sqrt{\varepsilon}/8$ 5.705E-02 5.111E-03 6.859E-04 1.982E-04
$\pi\sqrt{\varepsilon}/16$ 5.159E-02 3.669E-03 3.217E-04 4.312E-05
Table 1: Values of $M_{\text{LSA}}$ using neighbour strategy Q2P2
$dp$ $dq$ $2\sqrt{\varepsilon}$ $\sqrt{\varepsilon}$ $\sqrt{\varepsilon}/2$ $\sqrt{\varepsilon}/4$
$\pi\sqrt{\varepsilon}/2$ 3.447E-02 9.531E-03 4.683E-03 3.518E-03
$\pi\sqrt{\varepsilon}/4$ 2.722E-03 1.728E-04 4.318E-05 2.071E-05
$\pi\sqrt{\varepsilon}/8$ 6.586E-04 1.093E-05 6.700E-06 1.629E-06
$\pi\sqrt{\varepsilon}/16$ 3.819E-04 6.017E-06 5.660E-06 1.165E-06
Table 2: Values of $M_{\text{LSA}}$ using neighbour strategy Q4P2
$dp$ $dq$ $2\sqrt{\varepsilon}$ $\sqrt{\varepsilon}$ $\sqrt{\varepsilon}/2$ $\sqrt{\varepsilon}/4$
$\pi\sqrt{\varepsilon}/2$ 7.320E-02 2.183E-03 2.297E-04 7.624E-05
$\pi\sqrt{\varepsilon}/4$ 3.114E-02 4.129E-04 8.925E-06 4.336E-06
$\pi\sqrt{\varepsilon}/8$ 2.159E-02 1.653E-04 1.262E-05 7.480E-06
$\pi\sqrt{\varepsilon}/16$ 1.928E-02 1.130E-04 4.329E-06 2.171E-06
Table 3: Values of $M_{\text{LSA}}$ using neighbour strategy Q2P4
$dp$ $dq$ $2\sqrt{\varepsilon}$ $\sqrt{\varepsilon}$ $\sqrt{\varepsilon}/2$ $\sqrt{\varepsilon}/4$
$\pi\sqrt{\varepsilon}/2$ 2.054E-03 1.326E-05 2.365E-05 1.178E-05
$\pi\sqrt{\varepsilon}/4$ 6.237E-05 1.385E-05 1.663E-06 3.004E-06
$\pi\sqrt{\varepsilon}/8$ 1.719E-05 4.352E-06 1.125E-06 7.155E-07
$\pi\sqrt{\varepsilon}/16$ 1.596E-04 1.734E-05 1.428E-06 2.181E-07
Table 4: Values of $M_{\text{LSA}}$ using neighbour strategy Q4P4
3.2 Improved versions of FGGC
In Step 3 of algorithm 1, performing the least squares approximation for each wave-packet involves computing the pseudo-inserve $\bm{A}^{\dagger}$ and the matrix-vector multiplication $\bm{A}^{\dagger}\bm{f}$. This requires $O(n^{3}+n^{2})$ operations for each wave-packet. However, recalling eq. 3.9 and eq. 3.13 from the proof of theorem 2, we see that approximating $G(\bm{x};\bm{Q},\bm{P})$ using $G\left(\bm{x};\bm{\mathcal{N}}(\bm{Q},\bm{P};k)\right)$ is almost equivalent to approximating $G\left(\bm{x};\operatorname{frac}\bm{Q},\operatorname{frac}\bm{P}\right)$ using $G\left(\bm{x};\bm{\mathcal{N}}(\bm{0},\bm{0};k)\right)$, differing only by a constant factor in the approximation coefficients. It should be noted that, as the second equation in eq. 3.9 shows, $\widetilde{A}_{jk}$ is fixed with given mesh strategy and neighbour strategy, so $\widetilde{\bm{A}}^{\dagger}$ can be pre-computed independently. In doing so, to obtain the LSA coefficients for each off-grid wave-packet, it suffices to calculate $\widetilde{\bm{f}}$ and $\widetilde{\bm{c}}=\widetilde{\bm{A}}^{\dagger}\widetilde{\bm{f}}$, and then eq. 3.13 yields that the LSA coefficients $\left\{c^{(k)}\right\}_{k=1}^{n}$ can be obtained from $\left\{\widetilde{c}^{(k)}\right\}_{k=1}^{n}$.
By pre-computing $\widetilde{\bm{A}}^{\dagger}$, the LSA process requires only $O\left(n^{3}\right)+O\left(M\cdot n^{2}\right)$ operations instead of $O\left(M\cdot n^{3}\right)+O\left(M\cdot n^{2}\right)$ and reduces the computational cost. Consequently, we propose an improved version of the FGGC algorithm, presented in algorithm 2.
Algorithm 2 Frozen Gaussian grid-point correction (FGGC) - An improved version
  Step 0: Preparation
  Determine the mesh strategy $\Delta x=O(\varepsilon),\Delta q=O\left(\sqrt{\varepsilon}\right),\Delta p=O\left(\sqrt{\varepsilon}\right)$.
  Determine the neighbour strategy, including $n$ and $\bm{\delta q}^{(k)}$, $\bm{\delta p}^{(k)}$.
  Pre-compute $\widetilde{\bm{A}}^{\dagger}$.
  Step 1: Initial decomposition
  For each $\bm{q}$, compute $A(0,\bm{q},\bm{p})$ by eq. 2.7.
  Step 2: Time evolution
  For each wave-packet, evolve the ODE system eq. 2.3$\simeq$. 2.6 up to time $t$, and obtain the FGA variables $A,S,\bm{Q},\bm{P}$.
  Step 3: Least squares approximation
  For each wave-packet, calculate $\widetilde{\bm{f}}$ as eq. 3.9, let $\widetilde{\bm{c}}=\widetilde{\bm{A}}^{\dagger}\widetilde{\bm{f}}$, and then calculate $\bm{c}$ as eq. 3.13.
  Step 4: Wave reconstruction
  Compute the wave function based on eq. 2.24 and eq. 2.25.
Furthermore, motivated by the following two observations, we propose a multi-step version of the FGGC algorithm:
(1) Most of the computation time is concentrated in the ODE evolution step when the final time is large enough or $\varepsilon$ is not small enough;
(2) The FGGC evolves a set of on-grid wave-packets to another set of on-grid wave-packets. This process can be viewed as a linear map and can be learned by the solver.
Specifically, suppose $T_{\text{final}}=T_{\text{multi}}\cdot T_{\text{evo}}$, which means running the algorithm $T_{\text{multi}}$ times and each time evolving the wave function up to time $T_{\text{evo}}$. For a fixed time $T_{\text{evo}}$, a single on-grid wave-packet before evolution corresponds to $n$ on-grid wave-packets after evolution, each carrying a weight. This process can be viewed as a linear map, which takes an initial position as input and produces $n$ final positions and $n$ coefficients as output. For simplicity, the $n$ final positions can also be represented by a single final position once the neighbour strategy is determined. At each iteration, the algorithm begins with the initial decomposition. Then, the solver learns the evolution and splitting results consisting of $n+2$ triples (initial position, final position, and $n$ splitting coefficients) and improves efficiency in the subsequent runs. Once all the the wave-packets have been computed, the wave function can be reconstructed, completing one iteration of the algorithm. Repetition of the above procedure $T_{\text{multi}}$ times leads to the final state of the wave function. The multi-step version of FGGC is detailed in algorithm 3.
Algorithm 3 Frozen Gaussian grid-point correction (FGGC) - A multi-step version
  Step 0: Preparation
  Determine the mesh strategy $\Delta x=O(\varepsilon),\Delta q=O\left(\sqrt{\varepsilon}\right),\Delta p=O\left(\sqrt{\varepsilon}\right)$.
  Determine the neighbour strategy, including $n$ and $\bm{\delta q}^{(k)}$, $\bm{\delta p}^{(k)}$.
  Pre-compute $\widetilde{\bm{A}}^{\dagger}$.
  for $N_{\text{T}}=1,\dots,T_{\text{multi}}$ do
     Step 1: Initial decomposition
     For each $\bm{q}$, compute $A(0,\bm{q},\bm{p})$ by eq. 2.7, in which $u(0,\bm{x})$ should be substituted by $u((N_{\text{T}}-1)\cdot T_{\text{evo}},\bm{x})$.
     Step 2: Time evolution and least squares approximation
     For each wave-packet not learned yet: evolve the ODE system eq. 2.3$\simeq$. 2.6 up to time $T_{\text{evo}}$, calculate the approximation coefficients $\bm{c}$ using the pre-computed $\widetilde{\bm{A}}^{\dagger}$. The relationship between the initial position, final position and splitting coefficients is learned during the process.
     Step 3: Wave reconstruction
     Compute the wave function $u(N_{\text{T}}\cdot T_{\text{evo}},\bm{x})$ based on eq. 2.24 and eq. 2.25.
  end for
The advantages of this multi-step solver are: (1) One may only need to evolve a larger number of points during the first run of the algorithm, and subsequently evolve much fewer points in later runs, which may improve efficiency compared to evolve all points up to the final time. (2) The solver may eventually learn the entire map of the grid-points as the algorithm runs sufficiently many times, by then there would be no need to calculate the ODE evolution and LSA anymore. (3) More information about the wave function at intermediate stages can be obtained. However, this design also has its disadvantages: (1) The error accumulates with each run. (2) The algorithm is more efficient only if the ODE evolution step dominates the entire computation time of the algorithm.
4 Numerical Examples
In this section, we present several numerical examples to verify the accuracy and efficiency of the algorithm. For fixed mesh strategy, we aim to:
(1) Demonstrate that the error introduced by the least squares approximation is irrelevant to $\varepsilon$ numerically, by calculating $\left\|u_{\textup{FGA}}-u_{\textup{FGGC}}\right\|_{L^{2}}$ and observing its trend with respect to $\varepsilon$.
(2) Observe the improvement in efficiency, compared with FGA and TSSP.
(3) Study how different neighbour strategies affect the FGGC algorithm.
All our numerical experiments are conducted by CPU-parallel implementation (CPU-sequential implementation for the ones in the sequential scenario), on a system equipped with an 96-cores AMD Ryzen Threadripper 7995WX processor (384MB of L3 cache), 256GB of RAM and running Ubuntu 24.04. CPU-parallel implementations are compiled using the Intel IFX fortran compiler 2025.0.4, and parallel execution was facilitated by OpenMP. The CPU runtimes in the following subsections are all reported in seconds.
4.1 One-dimensional case
Example 1.
Take
$u_{0}(x)=\sqrt[4]{\frac{64}{\pi}}\operatorname{exp}\left(-32x^{2}+\frac{\mathrm{i}}{\varepsilon}x\right)\text{\ \ and\ \ }V(x)=1-\cos(\pi x).$ (4.1)
We solve in the spatial interval $[-2,2]$ with the final time $T=0.8s$.
We compute the FGGC solutions, FGA solutions and TSSP solutions, with the latter two as reference solutions. We take $\Delta x=\varepsilon$, $\Delta q=\sqrt{\varepsilon}/2$, $\Delta p=\pi\sqrt{\varepsilon}/8$ for the meshes of FGGC and FGA, and the time step $\Delta t=10^{-4}$ to solve the ODE system by the 4-th order Runge-Kutta method. The TSSP solution is computed with the mesh size of $\Delta x=\varepsilon$ and the time step $\Delta t=10^{-4}$.
The overall $L^{2}$ errors are shown in table 8 for the FGA and FGGC solutions with different neighbour strategies, where the reference solutions are computed using TSSP. It can be observed that FGGC has almost the same overall error as FGA. The error introduced by the LSA is shown in table 8, where the reference solutions are computed using FGA. It can be seen that the LSA error is sufficiently small, and decreases rapidly as the neighbour strategy being refined. Furthermore, we explore how the error varies with changes in $\varepsilon$. table 8 indicates that: the LSA error of the FGGC remains virtually unchanged with change in $\varepsilon$, which aligns with our conclusion in section 3.1.
Note that the FGA and the FGGC are highly parallelizable. Although TSSP can also be computed in parallel, achieving the expected level of parallelism may be hindered by the “butterfly operations” of the FFT algorithm. We present the runtimes of the three algorithms in both sequential and parallel scenarios as table 8 and table 8 respectively. For large $\varepsilon$, there is almost no significant improvement in the runtimes of FGGC compared to FGA. This is because the majority of the runtime is consumed by solving the ODE system, which makes the reconstruction step nearly imperceptible. However, as $\varepsilon$ decreases, the efficiency advantage of FGGC becomes increasingly pronounced. In comparison with TSSP, FGA and FGGC demonstrate greater efficiency, and the advantage is further enhanced in parallel computing scenarios.
4.2 Two-dimensional case
Example 2.
Take
$u_{0}(x,y)=\sqrt{\frac{64}{\pi}}\operatorname{exp}\left(-32(x^{2}+y^{2})+\frac{\mathrm{i}}{\varepsilon}(x+y)\right),$ (4.2)
and
$V_{1}(x,y)=1-e^{x^{2}+y^{2}}\text{\ \ or\ \ }V_{2}(x,y)=\frac{x^{2}+y^{2}}{2}.$ (4.3)
We solve in the spatial interval $[-2,2]^{2}$ with the final time $T=0.8s$.
Similarly to example 1, we take $\Delta x=\varepsilon$, $\Delta q=\sqrt{\varepsilon}/2$, $\Delta p=\pi\sqrt{\varepsilon}/8$, $\Delta t=10^{-4}$ for the mesh steps of FGGC and FGA, and take $\Delta x=\varepsilon$ and $\Delta t=10^{-4}$ for TSSP. The overall $L^{2}$ errors of the FGA solutions and the FGGC solutions are presented in table 13 and table 13. The errors introduced by the LSA are presented in table 13 and table 13. The results in table 13 and table 13 show that the LSA error is also sufficiently small, even compared to the discretization error or the truncation error. Based on the results of the four tables above, we once again conclude that the error introduced by the LSA can be sufficiently small, irrelevant to $\varepsilon$, and decreases rapidly as the number of neighbour points increases.
In the two-dimensional case we always perform the algorithm in the parallel scenario. We choose the case of potential $V_{1}$ to record the runtimes. The runtimes of the three algorithms are presented in table 13. The most significant discovery is that the FGGC saves a huge amount of computational resources in the reconstruction step for two-dimensional examples. The reconstruction step can be several dozen times faster than that of FGA. Although the least squares approximation may introduce additional computational cost, it is negligible compared to the huge time savings achieved in the reconstruction step. Furthermore, when $\varepsilon$ is sufficiently small, the TSSP requires an exceptionally long time. For instance, it takes approximately 3 days to obtain the TSSP solution when $\varepsilon=2^{-14}$, which we decided to pass this run in the experiments. In comparison, the highly optimized and parallelized FGGC algorithm can obtain a reliable result in less than 8 minutes.
4.3 Three-dimensional case
Example 3.
Take
$u_{0}(x,y,z)=\left(\frac{64}{\pi}\right)^{\frac{3}{4}}\operatorname{exp}\left(-32(x^{2}+y^{2}+z^{2})+\frac{\mathrm{i}}{\varepsilon}(x+y+z)\right),$ (4.4)
and
$V(x,y,z)=\frac{x^{2}+y^{2}+z^{2}}{2}.$ (4.5)
We solve in the spatial interval $[-2,2]^{3}$ with the final time $T=0.8s$.
In three-dimensional cases, the mesh steps of FGGC and FGA are set as follows: Take $\Delta x=\varepsilon$, $\Delta t=10^{-4}$. For $\varepsilon=1/2^{5}$ or $1/2^{7}$, take $\Delta q=\sqrt{\varepsilon/2}/2$ and $\Delta p=\pi\sqrt{\varepsilon/2}/8$. For $\varepsilon=1/2^{6}$ or $1/2^{8}$, take $\Delta q=\sqrt{\varepsilon}/2$ and $\Delta p=\pi\sqrt{\varepsilon}/8$. The TSSP method takes $\Delta x=\varepsilon$ and $\Delta t=10^{-4}$. The overall $L^{2}$ errors of the FGA solutions and the FGGC solutions are presented in table 16. The LSA errors of the FGGC solutions are presented in table 16. It can be seen that the error introduced by the LSA can be sufficiently small, irrelevant to $\varepsilon$, and decreases rapidly as the number of neighbour points increases.
Next, we analyze the efficiency of these three algorithms in parallel scenario. We present the runtimes of the three algorithms in table 16. Once again we observe that the FGGC saves a huge amount of computational resources in the reconstruction step. In the three-dimensional examples, solving the Schrödinger equation becomes extremely difficult. Both TSSP and FGA require an intolerably long computation time when $\varepsilon$ is slightly larger. However, thanks to the improvement in reconstruction efficiency, the FGGC can still complete the computation in a relatively short amount of time.
4.4 Multi-step FGGC Algorithm
Example 4.
Take
$u_{0}(x)=\sqrt[4]{\frac{64}{\pi}}\operatorname{exp}\left(-32x^{2}+\frac{\mathrm{i}}{\varepsilon}x\right),$ (4.6)
and
$V_{1}(x)=1-\cos(\pi x)\text{\ \ or\ \ }V_{2}(x)=\frac{x^{2}}{2}$ (4.7)
We solve in the spatial interval $[-2,2]$ with $\varepsilon=1/2^{12}$ and the final time $T=0.8s$, using the multi-step version FGGC.
In this subsection, we study the error performance and efficiency of algorithm 3. In example 4, we take the mesh steps as in section 4.1 and the neighbour strategy Q2P2. It should be noted that the two potential functions in this example both form a potential well, so the wave-packets remain confined within a small region during their evolution. With $T=0.8s$ and $k=10^{-4}$, the ODE system requires $8000$ steps for evolution. Therefore, we decompose the number of evolution steps $8000$ into several scenarios: $4000\times 2$, $2000\times 4$, $1000\times 8$, $500\times 16$ and $250\times 32$. The overall errors and runtimes for each of these decompositions are presented in table 22, table 22, and table 22.
First, we found that the multi-step FGGC further improves the efficiency for certain decompositions of evolution steps. It can be observed that the $1000\times 8$ decomposition achieves the highest efficiency. As the final time is divided into progressively smaller segments (i.e. $T_{\text{multi}}$ increases), the overall CPU runtime exhibits a pattern of initially decreasing and then increasing. This is because the number of steps in ODE evolution may decrease, while the number of times for initial value decomposition and reconstruction will increase with $T_{\text{multi}}$ increasing. Then we analyze the error performance. Intuitively, as the final time is divided into progressively tinier divisions, the accumulation of the iteration error also grows. The results in table 22 and table 22 reveal that the iteration error accumulates by a factor of a few to ten in our experiment, which is relatively acceptable.
Example 5.
Take
$u_{0}(x,y)=\sqrt{\frac{64}{\pi}}\operatorname{exp}\left(-32(x^{2}+y^{2})+\frac{\mathrm{i}}{\varepsilon}(x+y)\right),$ (4.8)
and
$V_{1}(x,y)=1-e^{x^{2}+y^{2}}\text{\ \ or\ \ }V_{2}(x,y)=\frac{x^{2}+y^{2}}{2}.$ (4.9)
We solve on the spatial interval $[-2,2]^{2}$ with $\varepsilon=1/2^{6}$ and final time $T=1.6s$, using the multi-step version FGGC.
Next, we proceed to solve the two-dimensional example. We also take the mesh steps as in section 4.2 and the neighbour strategy Q2P2 for example 5. We decompose the number of evolution steps $16000$ into the following scenarios: $8000\times 2$, $4000\times 4$, $2000\times 8$, $1000\times 16$ and $500\times 32$. The overall errors presented in table 22 and table 22 demonstrate that the multi-step algorithm does not introduce too much iteration error. The runtimes are presented in table 22. It can be observed that $2000\times 8$ decompostition achieves the best efficiency.
5 Conclusion and Discussion
In this paper, we introduce the frozen Gaussian grid-point correction for computing the solution of the linear Schrödinger equation in the semi-classical regime. This algorithm builds upon the frozen Gaussian approximation and employs a straightforward “on-grid correction” approach using the least squares approximation. The FGGC demonstrates significant potential for reducing the computational costs during the wave reconstruction step by decomposing those off-grid Gaussian wave-packets to on-grid counterparts, thereby enabling the use of highly efficient fast Fourier transform algorithm. We theoretically prove that the error for least squares approximation has a uniform upper bound independent of $\varepsilon$. Several numerical examples demonstrate that the error introduced by the least squares approximation is indeed irrelevant to $\varepsilon$ and decreases rapidly as more neighbours are added. The FGGC algorithm achieves significant efficiency improvements while introducing only a negligible error cost. Additionally, a multi-step version of the FGGC algorithm is also proposed, which further enhances computational efficiency. Finally, this algorithm can be readily extended to other multi-scale problems solvable via the FGA and significantly enhances the efficiency of solving such equations.